OpenStudy (anonymous):

What is the radius of the circle shown below? x^2+y^2-8x-10y-8=0

OpenStudy (anonymous):

I've gotten 7 - is that correct?

OpenStudy (campbell_st):

you need to complete the square for both x and y \[x^2 - 8x + y^2 - 10y = 8\] (-8/2)^2 = 16 (-10/2)^2 = 25 add 16 and 25 to both sides \[x^2 - 8x + 16 + y^2 - 10y + 25 = 8 + 16 + 25\] factorise \[(x -4)^2 +(y - 5)^2 = 7^2\] the radius is 7

OpenStudy (anonymous):

Thanks!

OpenStudy (campbell_st):

hope it helps